# Intensity-Efficient and Intensity-Positional Allocations

### 01 Mar 2026 v6

@@ 6 Eliciting Ordinal Intensities and Improving Upon Pareto-Efficient Allocations
For the “if” part, let $x$ be an arbitrary allocation that satisfies (i) and (ii). Assume to the contrary that $x$ is not Pareto efficient. Let $y$ be another allocation that Pareto-dominates $x$. This implies $s_{l}(y_{l},x_{l})\geq 0$ for every $l\leq n$, with strict inequality for some $l$. Let $N^{*}\subsetneq N$ consist of every agent $l$ for whom $s_{l}(y_{l},x_{l})=0$. By ($\mathbf{5}_{D}$), $s_{l}(y_{l},x_{l})=0\Leftrightarrow y_{l}=x_{l}$. Suppose $N^{*}=\emptyset$. From the above, this implies $s_{l}(y_{l},x_{l})>0$ for all $l\leq n$, with every agent assigned a different item under $x$ and $y$. If the envy-free set $N^{1}$ is non-empty, then this readily contradicts (i) for every $v_{l}\in N^{1}$. If $N^{1}=\emptyset$, then there is at least one $i>1$ such that the top-preference-coincident set $N^{i>1}$ is non-empty. Since $|V|=|A|=n$ holds by assumption, and recalling the notation that precedes the statement of Lemma 3, it follows from (ii) that for the specific agent $v_{1}^{i}\in N^{i>1}$ we must have $s_{1}^{i}(x_{1}^{i},x^{\prime}_{i})>0$ for all $x^{\prime}_{i}\in A\setminus\{x_{1}^{i}\}$, a contradiction. Now suppose $N^{*}\neq\emptyset$. Since $y$ Pareto-dominates $x$, the set $N^{**}:=N\setminus N^{*}$ that comprises every agent $l$ for whom $s_{l}(y_{l},x_{l})>0$ holds is non-empty. From the preceding argument we know that for every $v_{l}\in N^{**}$ it must be that $v_{l}\not\in N^{1}$ and $v_{l}\neq v_{1}^{i}\in N^{i>1}$, for all $i>1$. By $s_{l}(y_{l},x_{l})>0$ and $|A|=|N|$, there is some agent $v_{k}\neq v_{l}$ such that $y_{l}=x_{k}$. By ($\mathbf{5}_{D}$) and Pareto-dominance, $s_{k}(y_{k},x_{k})>0$. Since both $v_{k}$ and $v_{l}$ are improved upon by $y$, it follows that $v_{l}\in N^{i>1}$ and $v_{k}\in V^{j>1}$ for distinct top-preference-coincident sets $N^{i>1}$ and $N^{j>1}$ with respect to $A^{i>1}$ and $A^{j>1}$. By the homeo-monotonicity assumption, $A^{i}\cap A^{j}=\emptyset$. Since, by the definition of top-preference-coincident sets, $s_{l}(a,a^{\prime})>0$ and $s_{k}(b,b^{\prime})>0$ holds for all $a\in A^{i>1},a^{\prime}\in A\setminus A^{i>1}$ and $b\in A^{j>1},b^{\prime}\in A\setminus A^{j>1}$, and since, by (ii), it holds that $x_{l}\in A^{i>1}$ and $x_{k}\equiv y_{l}\in A^{j>1}$, the postulate $s_{l}(y_{l},x_{l})>0$ leads to a contradiction. Therefore, $x$ is Pareto-efficient.
@@
By definition,
$\begin{array}[]{ccccc}r_{l}(a_{i})\equiv\max\limits_{a_{k}\in A}s_{l}(a_{i},a_{k})&\geq&r_{l}(a_{j})\equiv\max\limits_{a_{k}\in A}s_{l}(a_{j},a_{k})&\Longleftrightarrow\\ s_{l}(a_{i},a_{k})&\geq&s_{l}(a_{j},a_{k})\quad\forall\;a_{k}\in A&\Longleftrightarrow\\ (a_{i},a_{k})&\dot{\succsim}_{l}&(a_{j},a_{k})\quad\forall\;a_{k}\in A&\Longleftrightarrow&\\ (a_{i},a_{j})&\dot{\succsim}_{l}&(a_{j},a_{i})&\Longleftrightarrow&a_{i}\succsim_{l}a_{j},\end{array}$
where the penultimate equivalence immediately follows from $(\mathbf{4_{\mathit{D}}})$, which in turn is implied by the postulate that $s_{l}$ represents $\dot{\succsim}_{l}$ in profile $S\in\mathcal{\widetilde{S}}$. ∎
@@
Since there are $n$ alternatives in $A$, the above shows that there are n1n-1 intensity-equivalence classes $a_{i},a_{j}]\in(A\times A)_{\dot{\sim}_{l}}$ such that $a_{i}\succ_{l}a_{j}$. By ($1$), $r_{l}(a_{i})\equiv\max\limits_{a_{k}\in A}s_{l}(a_{i},a_{k})$. This and the above-postulated $\succ_{l}$ ordering on AA jointly imply
@@
Proof of Proposition 3. By the definitions of $b_{l}$ and $r_{l}$ in (10) and (11), for any $S\in\mathcal{\widehat{S}}$ and every $l\leq n$ there is a strictly increasing function $f_{l}:\{0,1,2,\ldots,n-1\}\rightarrow\{0,1,2,\ldots,{{n}\choose{2}}\}$ such that, for all $a\in A$,
$r_{l}^{S}(a)$ $\equiv$ $f_{l}(b_{l}^{P_{S}}(a))$
Hence, for $x,y\in\mathcal{A}$ it holds that
$\sum_{i=1}^{n}b_{l}^{P_{S}}(x_{i})>\sum_{i=1}^{n}b_{l}^{P_{S}}(y_{i})$ $\Longrightarrow$ $\sum_{i=1}^{n}r_{l}^{S}(x_{i})>\sum_{i=1}^{n}r_{l}^{S}(y_{i})$ (39)
To prove the first claim, suppose to the contrary that $x\in\mathsf{IP}(S)$ but $x\not\in\mathsf{Bo}(S)$. Then, there must be $y\in\mathcal{A}$ such that $\sum\limits_{i=1}^{n}b_{l}^{P_{S}}(y_{i})>\sum\limits_{i=1}^{n}b_{l}^{P_{S}}(x_{i})$. By (39), this contradicts the postulate $x\in\mathsf{IP}(S)$. The second claim is proved in the main text. $\square$
Proof of Proposition 4. In the main text.
Proof of Proposition 5. It suffices to note that (4) and $(\mathbf{5}_{D})$ together imply that any $\dot{\succsim}_{l}\in\mathcal$$\widehat${I}} contains $|\dot{\succsim}_{l}|=2{n\choose 2}+n=(n-1)(n+1)$ comparisons between pairs in $A\times A$. $\square$
Proof of Proposition 6. Let $x$ be the Pareto-efficient allocation that is found in Step 2 under the random priority ordering $\gg$ drawn in the algorithm’s first step. In the worst-case scenario, Step 4 always—with the possible exception of the agent ranked second-last by $\gg$—performs exhaustive search and removes one agent and one object in every iteration: Starting with $v_{1},x_{1}$, after n1n-1 searches it removes $v_{1},x_{1}$. Continuing with $v_{2},x_{2}$, after n2n-2 searches it removes $v_{2},x_{2}$. $\cdots$ Continuing with $v_{n-1},x_{n-1}$, after one search it removes both $v_{n-1},x_{n-1}$ and $v_{n},x_{n}$, either by implementing the swap $x_{n-1}\leftrightarrow x_{n}$ or otherwise. Worst-case performance is therefore associated with a total number of searches equalling
$\begin{array}[]{llclcllll}\sum\limits_{m=1}^{n-1}(n-m)&=&n^{2}-n-\sum\limits_{m=1}^{n-1}m&=&n^{2}-\sum\limits_{m=1}^{n}m&=&&&\\ &&&&&&\\ &&n^{2}-\dfrac{n}{2}(n+1)&=&\dfrac{1}{2}n(n-1)&=&\displaystyle{{n\choose k}}&&\end{array}$
Given that a Pareto-efficient allocation can be found by RP in Step 2 in $O(n)$ time (Manlove, 2013, Section 6.2), it follows that the algorithm terminates in at most ${n}\choose{2}$} steps, hence in $O(n^{2})$ time.
It remains to be shown that, if the final allocation is $y\neq x$, then $yDx$. Suppose $n=3$ first. Write $x:=(x_{1},x_{2},x_{3})$. By the definition of Step 4, the transition from $x$ to $y$ involves a single pairwise swap. Thus, $y$ must be one of the following: (i) $(x_{1},x_{3},x_{2})$, implying $s_{2}(x_{3},x_{2})>s_{3}(x_{3},x_{2})>0$ or $s_{3}(x_{2},x_{3})>s_{2}(x_{2},x_{3})>0$; (ii) $(x_{3},x_{2},x_{1})$, implying $s_{1}(x_{3},x_{1})>s_{3}(x_{3},x_{1})>0$ or $s_{3}(x_{1},x_{3})>s_{1}(x_{1},x_{3})>0$; (iii) $(x_{2},x_{1},x_{3})$, implying $s_{2}(x_{1},x_{2})>s_{1}(x_{1},x_{2})>0$ or $s_{1}(x_{2},x_{1})>s_{2}(x_{2},x_{1})>0$. In each case, $yDx$.
Now suppose $n>3$ and $S\in\mathcal{\widehat{S}}$ is effectively homeo-monotonic. Denote by $(N^{i})_{i=1}^{k}$ and $(A^{i})_{i=1}^{k}$ the respective partitions of $N$ and $A$ into a (possibly empty) top-preference diverse set $N^{1}$ with respect to $A^{1}$ and top-preference-coincident sets $N^{i>1}$ with respect to $A^{i>1}$, $i\leq k$. By the properties of these sets and profile, and by the fourth step of the algorithm, irrespective of the $\gg$-order on $N$ there can be no swaps between $x_{i}$ and $x_{j}$ for $v_{i},v_{j}\in N^{1}$. Thus, $y_{i}=x_{i}$ for all $v_{i}\in N^{1}$. By the properties of $(N^{i>1},A^{i>1})$, moreover and the instruction in Step 4, there can be no swaps between $x_{i}$ and $x_{j}$ for $v_{i}\in N^{l>1}$ and $v_{j}\in N^{m>1}$ when $m,l\in\{2,\ldots,k\}$ are distinct. It follows that $y$ differs from $x$ by at most $n\over 2$} or $n-1\over 2$} pairwise swaps/transpositions if $n$ is even or odd, respectively, with each such $x_{i}\leftrightarrow x_{j}$ swap occurring within the same top-preference-coincident pair $(N^{l>1},A^{l>1})$. By Step 4, finally, every such swap reassigns the—by Pareto-efficiency of xx—jointly preferred item within the relevant pair in the direction of intensity-dominance. Hence, $yDx$. $\square$

### 14 Apr 2026 v8

@@ 6 Concluding Remarks
For the “if” part, let $x$ be an arbitrary allocation that satisfies (i) and (ii). Assume to the contrary that $x$ is not Pareto efficient. Let $y$ be another allocation that Pareto-dominates $x$. This implies $s_{l}(y_{l},x_{l})\geq 0$ for every $l\leq n$, with strict inequality for some $l$. Let $N^{*}\subsetneq N$ consist of every agent $l$ for whom $s_{l}(y_{l},x_{l})=0$. By Strictness, $s_{l}(y_{l},x_{l})=0\Leftrightarrow y_{l}=x_{l}$. Suppose $N^{*}=\emptyset$. From the above, this implies $s_{l}(y_{l},x_{l})>0$ for all $l\leq n$, with every agent assigned a different item under $x$ and $y$. If the envy-free set $N^{1}$ is non-empty, then this readily contradicts (i) for every $v_{l}\in N^{1}$. If $N^{1}=\emptyset$, then there is at least one $i>1$ such that the top-preference-coincident set $N^{i>1}$ is non-empty. Since $|V|=|A|=n$ holds by assumption, and recalling the notation that precedes the statement of Lemma 3, it follows from (ii) that for the specific agent $v_{1}^{i}\in N^{i>1}$ we must have $s_{1}^{i}(x_{1}^{i},x^{\prime}_{i})>0$ for all $x^{\prime}_{i}\in A\setminus\{x_{1}^{i}\}$, a contradiction. Now suppose $N^{*}\neq\emptyset$. Since $y$ Pareto-dominates $x$, the set $N^{**}:=N\setminus N^{*}$ that comprises every agent $l$ for whom $s_{l}(y_{l},x_{l})>0$ holds is non-empty. From the preceding argument we know that for every $v_{l}\in N^{**}$ it must be that $v_{l}\not\in N^{1}$ and $v_{l}\neq v_{1}^{i}\in N^{i>1}$, for all $i>1$. By $s_{l}(y_{l},x_{l})>0$ and $|A|=|N|$, there is some agent $v_{k}\neq v_{l}$ such that $y_{l}=x_{k}$. By Strictness and Pareto-dominance, $s_{k}(y_{k},x_{k})>0$. Since both $v_{k}$ and $v_{l}$ are improved upon by $y$, it follows that $v_{l}\in N^{i>1}$ and $v_{k}\in V^{j>1}$ for distinct top-preference-coincident sets $N^{i>1}$ and $N^{j>1}$ with respect to $A^{i>1}$ and $A^{j>1}$. By the homeo-monotonicity assumption, $A^{i}\cap A^{j}=\emptyset$. Since, by the definition of top-preference-coincident sets, $s_{l}(a,a^{\prime})>0$ and $s_{k}(b,b^{\prime})>0$ holds for all $a\in A^{i>1},a^{\prime}\in A\setminus A^{i>1}$ and $b\in A^{j>1},b^{\prime}\in A\setminus A^{j>1}$, and since, by (ii), it holds that $x_{l}\in A^{i>1}$ and $x_{k}\equiv y_{l}\in A^{j>1}$, the postulate $s_{l}(y_{l},x_{l})>0$ leads to a contradiction. Therefore, $x$ is Pareto-efficient.
@@
Let $w^{l}\in A$ be any $\succsim_{l}$-minimal alternative, i.e. such that $a\succsim_{l}w^{l}$ for all $a\in A$. We first show that
$r_{l}(a)=\max_{b\in A}s_{l}(a,b)=s_{l}(a,w^{l})\qquad\text{for all }a\in A.$
Fix $a,b\in A$. If $a\succsim_{l}b$, then $s_{l}(a,b)\geq 0$. Since $w^{l}$ is $\succsim_{l}$-minimal, $b\succsim_{l}w^{l}$ is true, so $s_{l}(b,w^{l})\geq 0$ too. By lateral consistency,
$s_{l}(a,w^{l})\geq\max\{s_{l}(a,b),\,s_{l}(b,w^{l})\}\geq s_{l}(a,b).$
If, instead, $b\succ_{l}a$, then $s_{l}(a,b)<0$, while $a\succsim_{l}w^{l}$ implies $s_{l}(a,w^{l})\geq 0$. Hence, again $s_{l}(a,w^{l})\geq s_{l}(a,b)$. Since this holds for every $b\in A$, it follows that $s_{l}(a,w^{l})$ is the maximum value of $s_{l}(a,\cdot)$ on $A$, i.e. $r_{l}(a)=s_{l}(a,w^{l})$ for all $a\in A$. Now let $a_{i},a_{j}\in A$. Then
$r_{l}(a_{i})\geq r_{l}(a_{j})$ $\Longleftrightarrow$ $s_{l}(a_{i},w^{l})\geq s_{l}(a_{j},w^{l})$ (38)
$\Longleftrightarrow$ $(a_{i},w^{l})\,\dot{\succsim}_{l}\,(a_{j},w^{l}).$ (39)
We will show that $(a_{i},w^{l})\,\dot{\succsim}_{l}\,(a_{j},w^{l})$ $\Leftrightarrow$ $(a_{i},a_{j})\,\dot{\succsim}_{l}\,(a_{j},a_{i})$. Suppose $(a_{i},w^{l})\,\dot{\succsim}_{l}\,(a_{j},w^{l})$. This is equivalent to $s_{l}(a_{i},w^{l})\geq s_{l}(a_{j},w^{l})$. By the definition of $w^{l}$, moreover, $s_{l}(a_{j},w^{l})\geq 0$. Suppose to the contrary that $(a_{j},a_{i})\,\dot{\succ}_{l}\,(a_{i},a_{j})$. This is equivalent to $s_{l}(a_{j},a_{i})>0>s_{l}(a_{i},a_{j})$. By (6), $s_{l}(a_{j},a_{i})>0$ and $s_{l}(a_{i},w^{l})\geq 0$ implies $s_{l}(a_{j},w^{l})>s_{l}(a_{i},w^{l})$, which is equivalent to $(a_{j},w^{l})\,\dot{\succ}_{l}\,(a_{i},w^{l})$. This contradicts our original postulate. Conversely, suppose $(a_{i},a_{j})\,\dot{\succsim}_{l}\,(a_{j},a_{i})$, and assume to the contrary that $(a_{j},w^{l})\,\dot{\succ}_{l}\,(a_{i},w^{l})$; that is, $s_{l}(a_{i},a_{j})\geq s_{l}(a_{j},a_{i})$ and $s_{l}(a_{j},w^{l})>s_{l}(a_{i},w^{l})$. By (6) and a similar reasoning we are now led to the absurd conclusion that $s_{l}(a_{i},w^{l})>s_{l}(a_{j},w^{l})$. Now, by the definition of the preference relation induced by $\dot{\succsim}_{l}$, $(a_{i},a_{j})\,\dot{\succsim}_{l}\,(a_{j},a_{i})$ is equivalent to $a_{i}\succsim_{l}a_{j}$. Therefore, from (39) we get
$r_{l}(a_{i})\geq r_{l}(a_{j})\Longleftrightarrow a_{i}\succsim_{l}a_{j},$
so $r_{l}$ is an ordinal utility representation of the preference relation induced by $\dot{\succsim}_{l}$. ∎
@@
Since there are $n$ alternatives in $A$, the above shows that there are n1n-1 intensity-equivalence classes $a_{i},a_{j}]\in(A\times A)_{\dot{\sim}_{l}}$ such that $a_{i}\succ_{l}a_{j}$. By (11), $r_{l}(a_{i})\equiv\max\limits_{a_{k}\in A}s_{l}(a_{i},a_{k})$. This and the above-postulated $\succ_{l}$ ordering on AA jointly imply
@@
Proof of Proposition 3. Let $x$ be intensity-positional at $S\in\widehat{S}$. Suppose, toward a contradiction, that $x$ is not Pareto-efficient. Then there exists some $y\in\mathcal{A}$ that Pareto-dominates $x$ at the induced preference profile $P_{S}$. That is, $y_{i}\succsim_{i}x_{i}$ for all $i\leq n$, with $y_{j}\succ_{j}x_{j}$ for some $j\leq n$. For each agent $i\leq n$, the Borda score function $b_{i}^{P_{S}}$ is an ordinal utility representation of $\succ_{i}$, and so is the intensity-rank score $r_{i}^{S}$ (Lemma 5). Hence there exists a strictly increasing function $f_{i}:\{0,1,\dots,n-1\}\to\{0,1,\dots,{n\atopwithdelims( )2}\}$ such that $r_{i}^{S}(a)=f_{i}\bigl(b_{i}^{P_{S}}(a)\bigr)$ for all $a\in A$. Since $y_{i}\succsim_{i}x_{i}$ for all $i$ and $y_{j}\succ_{j}x_{j}$ for some $j$, it follows that $b_{i}^{P_{S}}(y_{i})\geq b_{i}^{P_{S}}(x_{i})$ for all $i$, with strict inequality for at least one $i$. Because each $f_{i}$ is strictly increasing, we obtain $r_{i}^{S}(y_{i})\geq r_{i}^{S}(x_{i})$ for all $i$, with strict inequality for at least one $i$. Therefore, $\sum\limits_{i=1}^{n}r_{i}^{S}(y_{i})>\sum\limits_{i=1}^{n}r_{i}^{S}(x_{i})$, contradicting the assumption that $x$ is intensity-positional at $S$. Hence, any intensity-positional allocation is Pareto-efficient. $\square$
Proof of Proposition 4. In the main text. $\square$